BACK STANDARD VTEW PRİNTER VERSİON P9.029 G0 Multipart Part 1 The cantilever beam shown is subjected to a concentrated load of P- 125 kN. The cross-sectional dimensions of the t ectangular tube shape are shewn, where b- 150 mm, d 250 mm -8 mm, and 60 mm. Determine (a) the stear stress at point K, located d- b) the maximum horizontal shear stress in the rectangular tube shape. which s loceted d - 60 mm below the cent of the rectanguilar tube shspe. d of the rectangular tube shape. УК : (typ.) Determine the moment of inertia I, for the cross-section abou: the z centroidal axis Answer: the tolerance is +/-2% Click if you would like to Show work for this question: asnstenwoch Attempts: Unlimited SAVE FOR LATER SUBHIT ANEIWER Part 2 connpute the val of of the first r ont of area Q t is asociated with Pol t κ which s located d 60 mm be ow the cent old of the rectangular tube shape. Answer: Q the tolerance is +/-2% Click if you would like to Show Work for this ouestinen
I STANDARD VIEW PRINTER VERSION Part 2 Compute the value of of the first moment of area Q that is associated with point, which is located Ơ-60 mm below the centroid of the rectangular tube shape. Answer: Q-mm2 the tolerance is +/-2% Click if you would like to Show Work for this question: Open Show Work Attempts: Unlimited SAVE FOR LATER SUBMIT ANSW Part 3 Determine the shear stress at point K. Answer: K- tho tolerance is +/-2% Click if you would like to Show Work for this question: Qoen Show Work MPa. Attempts: Unlimited SAVE FOR LATER SUBHIT ANSWER Part 4 Determine the maximum value of the first moment of area Q for the cross-section. Answer: the tolerance is +/-2% Click if you would like to Show Work for this question: Open Show Work
Part 5 Determine the maximum horizontal shear stress in the rectangular tube shape. Answer: max - the tolerance is +/-2% Click if you would like to Show Work for this question: MPa Open Show Work
Answer #1

Q1 part (a)

\tau=\frac{F*A*\widehat{y}}{I*b} (eq. 1)

where

\tau = Shear stress

F= Applied load

\widehat{y}= mid point of area of interest

I = Moment of inertia

A= area

b = width at point of interest

In our case

F= 125 kN

A(flange)*\widehat{y}=B(\frac{d-(d-2*t)}{2})*(\frac{d-(d-2*t)}{4}) (eq. 2)

A(web )*\widehat{y}=2*t(\frac{(d-2*t)}{2}-y)*\frac{1}{2}(\frac{(d-2*t)}{2}+y) (eq. 3)

A*\widehat{y} (total) =A(flange)*\widehat{y} + A(web )*\widehat{y}

= 85511.99 mm^3

  I =\frac{bd^3-hk^3}{12} (eq.4)

here h=b - 2*t=150-2*8 =134

k= d - 2*t=250-2*8 = 234

I = 52235072 mm^4

A= (b*d)-(h*k)/2-(2*t*y) = 2112 mm^2 (here y is 60 mm that is given in question.)

b = 2*t= 2*8= 16 mm

Putting these values in eq. 1

\tau= 12.789539  N/mm^2 at point 60 mm below centroid.

Q1 part(b)

In eq. 2 and eq . 3

take y= 0

as maximum stress occurs at centroid

then

\widehat{y}= 62.5 mm

A= 3072 mm^2

\tau= 28.71633  N/mm^2

Part 2

\widehat{y} (total) = d/2 -  A(flange)*\widehat{y} + A(web )*\widehat{y} /A

=250/2- 85511.99/3872

= 102.91 mm

Part 3

answered above at point k stress

Part 4

area calculation in part 1